A car travels at 40 km/h for 50 minutes, then at 90 km/h for 6.0 km. what is the average velocity?
1
Expert's answer
2012-12-20T09:22:29-0500
V1=40 km/h & T1=50 minutes=5/6 hours
V2=90 km/h & S2=6 km
V_av=S_total/T_total
S_total - total distance
T_total - total time
S1=T1*V1 - the distance covered in the first section
S_total=S1+S2=T1*V1+S2
T2=S2/T2 - the time spent on the second section
T_total=T1+S2/T2
V_av=(T1*V1+S2)/(T1+S2/T2)=(40*5/6+6)/(5/6+6/90)=1180/27=43.70 km/h